Find vectors u,v "\in" R2 such that u is a scalar multiple of (1,3), v is orthogonal to (1,3), and (1,2) = u +v.


1
Expert's answer
2021-07-14T12:49:23-0400

Let "u=(u_1, u_2), v=(v_1, v_2)."


"u" is a scalar multiple of "(1, 3):" "u_2 =3u_1."


"v" is orthogonal to "(1,3):v_1+3v_2=0=>v_1=-3v_2"


"u+v=(1,2):u_1+v_1=1, u_2+v_2=2"


Hence


"\begin{matrix}
 u_1-3v_2= 1\\
 3u_1+v_2 =2
\end{matrix}"

"\begin{matrix}
 u_1=3v_2+ 1\\
 9v_2+3+v_2 =2
\end{matrix}"


"\begin{matrix}
 u_1=0.7\\
 v_2 =-0.1
\end{matrix}"

Then


"u=(0.7, 2.1),"

"v=(0.3, -0.1)."